### v1

# Dual Magnetic and Electric Dipole Symmetry: Pseudo-Angular Momentum in Parity Space and the Electric Landé $g$-Factor

Abstract
Electric dipole moments (EDMs) are sensitive probes of fundamental symmetries and central to searches for physics beyond the Standard Model. We present a symmetry-based description of EDMs, analogous to the Zeeman effect, that places magnetic and electric dipoles on equal footing under electromagnetic duality. In the hydrogenic problem, a static $\vec{B}$ (pseudovector) couples to the angular-momentum operator $\hat{\vec{J}}$ and reduces $SO(4)$ to the axial $SO(2)$ generated by $\hat{J}_{z}$. In contrast, a static $\vec{E}$ (polar vector) couples within a fixed $n$ manifold to a scaled Runge-Lenz operator $\hat{\vec{A}}_{\rm sc}$, mixes opposite parities, and preserves $SO(2)\times SO(2)$ generated by $\hat{J}_{z}$ (axial) and $\hat{A}_{{\rm sc},z}$ (dynamical). This motivates a pseudo-angular-momentum operator $\hat{\vec{J}}_{p}$, built from $\hat{\vec{A}}_{\rm sc}$, and an electric Landé factor $g_{E}$, so that the orbital dipole operator reads $\hat{\vec{d}}_{\rm orb}=g_{E}\,d_{B}\,\hat{\vec{J}}_{p}/\hbar$, with the natural electric unit $d_{B}=ea_{0}=\frac{2\mu_{B}}{c\alpha}$ ($a_{0}$ and $\mu_{B}$ are the Bohr radius and magneton respectively). As an example, Stark mixing of $2s$ and $2p_{m=0}$ in hydrogen yields $|\langle d_{\rm orb}\rangle|=3d_{B}$ (i.e. $g_{E}=3$). Following Ohanian’s magnetisation formalism for the Zeeman effect, we construct its electric dual; the wavefunction’s microscopic polarisation, which has a non-zero curl defining an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$. The EDM expectation value becomes $\langle\hat{\vec{d}}_{tot}\rangle=-\frac{\epsilon_{0}}{2}\int\vec{r}\times\vec{J}_{m}\,d^{3}r=d_{B}\big[g_{E}\,\langle\hat{\vec{J}}_{p}\rangle/\hbar+g_{E}^{e}\,\langle\hat{\vec{S}}\rangle$$\hbar\big$], $g_{E}^{e}=\frac{2d_{\rm int}}{d_{B}}$. Here the spin operator, $\hat{\vec{S}}$, encodes any intrinsic EDM, $d_{\rm int}$, from symmetry-violating interactions, while $\hat{\vec{J}}_{p}$ captures the Stark-induced pseudo-angular momentum set by Runge-Lenz symmetry. This dual framework shows that induced EDMs arise from effective circulating probability magnetic currents, directly mirroring magnetic dipole moments arising from circulating probability electric currents.
I Introduction
Electric dipole moments (EDMs) are exceptionally sensitive probes of fundamental symmetries and play a central role in the search for physics beyond the Standard Model. In atomic and molecular systems, EDMs can arise through two distinct mechanisms: as intrinsic EDMs, aligned with the particle’s real spin and signaling new CP-violating physics, or as induced EDMs, generated by parity mixing of orbital states and naturally described by a pseudo-angular-momentum degree of freedom, as in the linear Stark effect. Although these mechanisms are conceptually different, one tied to spin in position space, the other to pseudo-angular-momentum in parity space, both manifest experimentally as level shifts under external electric fields, a signature exploited in high-precision spectroscopy and EDM searches.
The linear Stark effect is a well-known example of an EDM arising from orbital dynamics, and is the electric analogue of the Zeeman effect. This phenomenon describes the splitting or shifting of atomic energy levels in a static electric field [1]. For a nonzero first-order shift to occur, the system must contain states of opposite parity that are sufficiently degenerate. In excited hydrogen, for instance, the $n=2$ manifold includes the degenerate $2s$ and $2p$ orbitals. An applied electric field mixes these states, distorting the electron cloud and inducing a dipole moment aligned with the field, which breaks inversion symmetry and couples to the orbital motion, producing a first-order energy shift [2, 3].
In contrast, EDMs associated with intrinsic spin are expected to be permanent and arise from violations of discrete symmetries, particularly parity (P) and time reversal (T) invariance. Within quantum field theory, such EDMs emerge from CP-violating operators and are predicted to couple directly to the spin vector via terms such as,
$\mathcal{L}_{\text{EDM}}=-\frac{1}{2}d_{int}\bar{\psi}\sigma^{\mu\nu}\gamma^{5}\psi F_{\mu\nu}.$ (1)
Here, $d_{int}$ is the intrinsic EDM of the particle (in units of $C\cdot m$), $\psi$ is the Dirac spinor field, $\sigma^{\mu\nu}=\frac{i}{2}\gamma^{\mu},\gamma^{\nu}$ is the antisymmetric spin tensor, $F_{\mu\nu}$ is the electromagnetic field strength tensor and $\gamma^{5}$, ensures the interaction violates both parity (P) and time-reversal (T) symmetry. The detection of a spin-aligned EDM in a fundamental particle like the electron or neutron would constitute direct evidence of physics beyond the Standard Model and provide insights into the baryon asymmetry of the universe.
Despite decades of effort, no intrinsic EDM has been observed to date. Current experiments place stringent limits on possible CP-violating effects[4, 5, 6, 7, 8], driving continued interest in EDM searches across atomic, molecular, and solid-state platforms. Theoretical frameworks beyond the Standard Model such as supersymmetry, left-right symmetric models, and theories of baryogenesis often predict EDMs just below current experimental sensitivity.
Complementary to these quantum approaches, insights from condensed matter and electromagnetic energy generation provide an alternative view of EDM-like phenomena. For example, modern polarization theory models permanent electric polarization in materials using the Berry phase formalism, where each unit cell contributes a quantized dipole moment [9, 10, 11]. This macroscopic charge separation is analogous to intrinsic EDMs in fundamental systems. Active voltage sources also create permanent polarizations [12, 13, 14], and reflect symmetry-breaking mechanisms similar in form to those studied in particle physics[15, 16].
In this work, we build on the magnetic-electric analogy to develop a semi-classical model of the EDM, inspired by the definition of the Bohr magneton, and the analysis of Ohanian [17]. To establish the connection between the magnetic moment and the spin of the electron, Ohanian showed that spin can be viewed as arising from a circulating flow of energy in the wave field, analogous to the angular momentum carried by a classical circularly polarized wave. In doing so, he introduced an effective probability current density, similar to bound current in a macroscopic magnetized material, defined as the curl of an effective microscopic magnetization field constructed directly from the electron wavefunction.
By duality, we construct an electric analogue of the Bohr magneton, which we term the “Bohr EDM”, and introduce an effective probability magnetic current density, associated with the curl of an effective microscopic polarization field constructed directly from the electron wavefunction. In this framework, the induced EDM in a system, such as the Stark-mixed hydrogen doublet, is naturally described by a pseudo-angular momentum degree of freedom. Unlike “real” orbital angular momentum or spin, they do not represent a physical rotation in space, instead they reflect the extent of parity mixing between opposite-parity states and determine the structure of the induced polarization. This dual perspective offers a unifying foundation that links induced EDMs in atomic systems, intrinsic spin EDMs of fundamental particles, polarization phenomena in condensed matter physics, and active dipole sources that provide current and voltage, and radiate electromagnetic energy.
II Magnetic Dipole Moments
II.1 The Semi-Classical Bohr Magnetron
The Bohr magneton arises naturally from a semi-classical model of an electron orbiting a nucleus, as in the Bohr model of the atom. In this picture, an electron of charge e-e and mass $m_{e}$ moves in a circular orbit of radius $r$ with speed $v$, generating a current, $I$.
$I=\frac{e}{T}=\frac{ev}{2\pi r}$ (2)
To draw a classical parallel, consider a loop of current $I$ encircling an area $\vec{A}$. The magnetic dipole moment associated with this loop is given by:
$\vec{\mu}=I\vec{A}.$ (3)
For an electron in a circular orbit of radius $r$ and speed $v$, the current is $I=e/T=ev/(2\pi r)$, and the area is $A=\pi r^{2}$, yielding:
$\mu=\frac{evr}{2}=\frac{eL}{2m_{e}},$ (4)
where $L=m_{e}vr$ is the orbital angular momentum.
In the Bohr model, the quantisation condition imposes $L=\hbar$, for the lowest energy level, which yields a magnetic moment
$\mu_{B}=\frac{e\hbar}{2m_{e}},$ (5)
defining the Bohr magneton. This derivation illustrates how the calassical concept of a circulating charge loop, combined with quantum angular momentum quantisation, leads to a fundamental unit of magnetic moment in atomic physics.
II.2 Zeeman Effect due to Intrinsic and Orbital Magnetic Moments
In the presence of an external static and uniform magnetic field $\vec{B}$, the perturbation Hamiltonian arises from the interaction of the electron magnetic moments with $\vec{B}$.
$\hat{H}_{Z}=-\hat{\vec{\mu}}\cdot\vec{B}$ (6)
where the magnetic moment, $\vec{\mu}=\vec{\mu}_{L}+\vec{\mu}_{S}$, the sum of the orbital and intrinsic spin magnetic moments.
$\hat{\vec{\mu}}_{L}$ $=-\frac{\mu_{B}}{\hbar}\hat{\vec{L}}$, (7)
$\hat{\vec{\mu}}_{S}$ $=-g_{S}\frac{\mu_{B}}{\hbar}\hat{\vec{S}}$.
Here $g_{S}$ is the electron spin $g$-factor, which in Dirac theory equals 2, and in QED $\simeq 2.0023192$. While orbital moments follow from classical current loops, the spin moments are considered to have no classical counterpart.
This perturbation leads to lines splitting into several components in the presence of the magnetic field, with Zeeman corrections characterised by,
$\hat{H}_{M}=g_{M}\frac{\hat{J}_{z}}{\hbar}\mu_{B}B_{z};~\Delta E_{M}=g_{M}m_{j}\mu_{B}B_{z},$ (8)
where $\hat{H}_{M}$ is the the Zeeman Hamiltonian, $\hat{J}_{z}$ is the z-component of the total angular momentum, $\hat{\vec{J}}$, while, $\Delta E_{M}$ is the energy shift of each Zeeman sub-level, $g_{M}$ is the magnetic Landé g-factor, and $m_{j}=-J,-J+1...,J$, is the magnetic quantum number associated with the total angular momentum projection.
II.3 Ohanian’s Model of Spin: Magnetic Moment from Internal Magnetisation
In his seminal paper “What is Spin?” [17], Ohanian provides a physically motivated description of the electron’s magnetic moment by modeling it in terms of an effective probability quantum current distribution, despite the fact that spin has no classical analog in terms of rotating charge. His treatment shows that the electromagnetic properties of spin, specifically the magnetic moment, can be derived from a localised magnetisation density constructed from the Dirac wavefunction.
Ohanian introduces a microscopic magnetization vector field $\vec{M}$ defined as
$\vec{M}=\mu_{B}\psi^{\dagger}\vec{\Sigma}\psi,$ (9)
where $\psi$ is the Dirac spinor, and $\vec{\Sigma}$ represents the spin matrices. This magnetization leads to an effective bound probability current density,
$\vec{J}_{b}=\nabla\times\vec{M},$ (10)
analogous to classical magnetization currents. The magnetic moment $\vec{\mu}$ associated with this probability current is:
$\vec{\mu}=\frac{1}{2}\int\vec{r}\times\vec{J}_{b}\,d^{3}r.$ (11)
Using the relation to magnetization:
$\vec{\mu}=\int\vec{M}\,d^{3}r,$ (12)
Ohanian recovers the quantum result:
$\hat{\vec{\mu}}_{S}=-g_{S}\frac{\mu_{B}}{\hbar}\hat{\vec{S}},\quad g_{S}=2.$ (13)
Ohanian’s construction reveals that although spin cannot be interpreted as literal circulating charge, its associated magnetic moment behaves as if it originates from an effective probability current density. This provides a bridge between classical electromagnetism (magnetization and bound currents) and quantum field theory (spinor structure and symmetry). The result reinforces the idea that the magnetic moment of a spin-$\frac{1}{2}$ particle is not merely a formal operator feature, but corresponds to a real, observable electromagnetic quantity tied to internal degrees of freedom in the quantum field.
III Electric Dipole Moments
So far, no particle has been observed to have a nonzero intrinsic EDM. In the Standard Model, such EDMs are predicted to exist but at levels far below current experimental sensitivity (for example, the electron EDM is expected to be about $10^{-38}\,\mathrm{e\cdot cm}$, whereas the neutron is about $10^{-31}\,\mathrm{e\cdot cm}$). State-of-the-art experiments constrain the electron EDM to be smaller than $1.3\times 10^{-29}\,\mathrm{e\cdot cm}$ and the neutron EDM below $1.0\times 10^{-26}\,\mathrm{e\cdot cm}$, with similar null results for the muon and proton. These increasingly tight bounds are crucial because a confirmed intrinsic EDM would signal new CP-violating physics beyond the Standard Model, possibly helping to explain the matter-antimatter asymmetry of the Universe. By contrast, the electric dipole moments that appear in the linear Stark effect of hydrogen or polar molecules are not intrinsic but induced. For example, in hydrogen they arise from parity mixing of nearly degenerate orbitals, such as the $2s$ and $2p_{0}$ states. In this case, the dipole operator acts on a pseudo-angular momentum degree of freedom, defined in the two-level parity subspace, and the resulting EDM depends on the external field and the mixing angle.
III.1 Semi-Classical Model of a Quanta of Permanent Polarization and dipole moment
Figure 1: Semi-classical representations of the Bohr EDM, $\vec{d}_{B}$. (a) Ideal semi-classical unit cell of quantised constant permanent polarization, $\Delta\vec{P}=\frac{e}{\pi a_{e}^{2}}\hat{z}$. In this model, the total angular momentum, and net electrical and magnetic charge are zero. However, because the curl of the polarization is non-zero, an effective magnetic current appears on the radial surface, which can be modelled as a quanta of a south and north pole ($\mp q_{m}=\mp\frac{h}{e}$) rotating in opposite directions, and is related to pseudo-spin (or pseudo-angular momentum) in parity space rather than position space. (b) Classical Bohr EDM, represented by two point charges $\pm e$ separated by a distance of the Bohr radius. (c) Schematic representing a semi-classical intrinsic Bohr EDM, connected to the position space spin vector, $\vec{S}$, related to an effective magnetic current of $\vec{I}_{m}$ at a radial distance of $a_{0}$.
Similar to the semi-classical atomic Bohr magneton calculation one can use a similar technique to derive the semi-classical atomic dual electrical analogue. To do this we conceive a toy model model of two quanta of opposite signed magnetic charge $\pm\frac{h}{e}$ as shown in Fig.1, rotating at the the electron Bohr radius, $a_{0}=\frac{\hbar}{m_{\mathrm{e}}c\alpha}$, creating a EDM. Here $\alpha=\frac{e^{2}}{4\pi\varepsilon_{0}\hbar c}$ is the fine structure constant, $\epsilon_{0}$ the permittivity of free space, $c$ the speed of light.
In this model, the total net electrical and magnetic charge as well as angular momentum is zero. However a finite effective magnetic current exits, which we pose is not generated by mechanical orbital motion, but by pseudospin parity mixing. Given that $q_{m}=\frac{h}{e}$, the magnetic current at the boundary is given by,
$\vec{I}_{m}=-\frac{2q_{m}v}{2\pi a_{0}}\hat{\phi}=-\frac{e}{2\pi a_{0}^{2}m^{*}\epsilon_{0}\alpha c}(a_{0}m^{*}v)\hat{\phi}$ (14)
Here $m^{*}$ is the effective mass of the atomic particle related to the permanent dipole, which can vary from $0.01m_{e}$ to very large.
Assuming the total mass per unit cell (Fig.1(c)) is equivalent to the effective mass of the electron, then we assume $m^{*}\sim\frac{m_{e}}{2}$ (two magnetic charged particles), and equating $a_{0}m^{*}v\sim\hbar$, as in the derivation of the Bohr magneton in (5), then the magnetic current becomes, $\vec{I}_{m}=-\frac{e\hbar}{\pi a_{0}^{2}m_{e}\epsilon_{0}\alpha c}\hat{\phi}=-\frac{2\mu_{\mathrm{B}}}{\pi a_{0}^{2}\epsilon_{0}\alpha c}\hat{\phi}$. Thus, we can define the “Bohr EDM” in terms of the effective azimuthal magnetic current, via a left hand rule,
$d_{B}=-\epsilon_{0}I_{m}\pi a_{0}^{2}=\frac{2\mu_{\mathrm{B}}}{c\alpha}=ea_{0},$ (15)
the dual of the Bohr magneton in (5). This dipole moment is equivalent to two charges $\pm e$ separated by a distance of $a_{0}$ as shown in Fig. 1 (b), which also means $\vec{a}=a_{0}\hat{z}$, so Fig. 1(a) and 1(b) are equivalent.
In modern polarization theory, the bulk polarization is defined “modulo a quantum”, along with a Berry phase defined modulo $2\pi$, and is given by
$\Delta\vec{P}\;=\;\frac{e\,\vec{a}}{V},$ (16)
where $\vec{a}$ is any lattice vector and $V$ the unit-cell volume. This follows from an electron in a periodic potential, which gives the $U(1)$ gauge freedom (Bloch’s theorem); $\psi_{n\vec{k}}\!\to\!e^{-i\vec{k}\!\cdot\!\vec{a}}\psi_{n\vec{k}}$, which shifts the Berry phase by $2\pi$ and $\vec{P}$ by $e\vec{a}/V$. Here we consider a quasi-1D “cell” as shown in Fig.1 (c) of length $a$ (along $\hat{\vec{z}}$) and cross-section $A_{\!\mathrm{eff}}$, so the effective volume is, $V_{\!\mathrm{eff}}\;=\;A_{\!\mathrm{eff}}\,a$. In our construction $2\pi$ winding of the internal phase is equivalent to choosing $\vec{a}=a\hat{\vec{z}}$. Then from (16)
$\Delta P_{z}\;=\;\frac{e\,a}{V_{\!\mathrm{eff}}}\;=\;\frac{e}{A_{\!\mathrm{eff}}}\;=\;\frac{e}{\pi a_{0}^{2}}.$ (17)
This is equivalent to a single cell of a super cell in modern polarization theory [10], with a quanta of permanent polarization, $\Delta P_{z}$ of size on the order of the Bohr radius, where $A_{\!\mathrm{eff}}=\pi a_{0}^{2}$, as shown in Fig.1. Asserting that $\vec{a}=a_{0}\hat{z}$, the dipole moment can then be calculated from $\vec{d}_{B}=\int\Delta P_{z}d^{3}r=ea_{0}\hat{z}=d_{B}\hat{z}$, consistent with (15).
III.2 Dual Ohanian Model: Dipole moment from Internal Polarisation
III.2.1 Bohr Electric Dipole Moment from Spin
As discussed previously, In Ohanian’s seminal work, the intrinsic magnetic moment of the electron was interpreted semiclassically as arising from an effective circulating energy flow of the electron wavefunction, generating an angular momentum and associated magnetic dipole [17]. The effective magnetization density was defined from the bilinear spinor quantity $\psi^{\dagger}\vec{\Sigma}\psi$, leading to the construction of an effective probability current density $\vec{J}_{b}=\nabla\times\vec{M}$, from which the Bohr magneton, $\mu_{B}$, naturally arises.
In analogy, we pose a putative dual picture in which a permanent EDM arises from an internal effective probability magnetic current distribution, that creates a permanent polarization by a left hand rule. This provides a complementary viewpoint in which the EDM can be understood as the macroscopic manifestation of internal field circulations analogous to those producing magnetic dipole moments. Thus, by invoking an analogue dual construction, one may define a spin operator, $\hat{\vec{S}}$, associated with the following internal microscopic polarization density,
$\vec{P}(\vec{r})$ $=d_{B}\,\psi^{\dagger}(\vec{r})\,\vec{\Sigma}\,\psi(\vec{r}),\qquad d_{B}\equiv ea_{0},$ (18)
$=\frac{2d_{B}}{\hbar}\psi^{\dagger}(\vec{r})\hat{\vec{S}}\psi(\vec{r}),\qquad\hat{\vec{S}}\equiv\frac{\hbar}{2}\bm{\sigma}.$ (19)
where, $\bm{\sigma}$ denotes the vector of Pauli operators acting on the two-level parity space, $\vec{\Sigma}$ is the Pauli spin matrix vector and $\psi$ the Dirac or Pauli spinor. Adopting the dual (Weber) convention for magnetic current, the curl of $\vec{P}$ defines an effective probability magnetic current density,
$\vec{J}_{m}\;=\;-\,\frac{1}{\epsilon_{0}}\nabla\times\vec{P},$ (20)
which plays the role dual to Ohanian’s effective current from $\vec{M}$. Since, $\|\langle\vec{\Sigma}\rangle\|=1$, the expectation value of the electric dipole moment is then
$\langle\vec{d}\rangle=\int\vec{P}(\vec{r})\,d^{3}r\;=\;d_{B},$ (21)
consistent with
$\hat{\vec{d}}=\frac{g_{E}^{S}d_{B}}{\hbar}\hat{\vec{S}}~~g_{E}^{S}=2,$ (22)
where $g_{E}^{S}$ is the electric Landé g-factor. Thus, a fully polarised single-particle state carries the “Bohr” electric dipole moment $d_{B}=ea_{0}$. This gives rise to an effective electric dipole moment associated with the spin degree of freedom. The spin carries an intrinsic dual moment generated by the internal microscopic polarization of the wavefunction. This construction provides a dual Ohanian model that allows a semiclassical interpretation of an effective intrinsic electric dipole moment.
In this framework, the magnetic moment $\vec{\mu}$ due to a circulating electric current is mirrored by an electric dipole moment $\vec{d}$ due to a circulating magnetic current:
$\langle\vec{d}\rangle=\frac{\epsilon_{0}}{2}\int\vec{r}\times\vec{J}_{m}\,d^{3}r,$ (23)
with a semi-classical picture shown in Fig.1(c). This construction preserves the correct transformation properties under parity (P) and time-reversal (T), making it consistent with known symmetry-breaking signatures of EDMs. However, in reality no EDMs coupled to intrinsic spin have yet to been observed, and if they do exist are significantly smaller than Bohr EDM.
III.2.2 Intrinsic and Orbital Polarization
If the electron carries an intrinsic electric dipole moment, $d_{int}$, the local microscopic polarization density of the electron wavefunction would be of the form,
$\vec{P}_{\rm e}(\vec{r})=d_{int}\,\psi^{\dagger}(\vec{r})\,\bm{\Sigma}\,\psi(\vec{r})\;\;\approx\;\;\frac{2d_{int}}{\hbar}\,\vec{s}(\vec{r}),$ (24)
consistent with
$\hat{\vec{d}}_{e}=\frac{g_{E}^{e}d_{B}}{\hbar}\hat{\vec{S}}~~~~g_{E}^{e}=\frac{2d_{int}}{d_{B}},$ (25)
where $\vec{s}(\vec{r})=\psi^{\dagger}(\vec{r})\frac{\hbar}{2}\bm{\sigma}\psi(\vec{r})$ is the nonrelativistic spin density and $g_{E}^{e}$ is the effective electric Landé g-factor of the electron. To understand how the electron behaves in an electric field we consider the interaction term in (1). To interpret this interaction physically, we examine its nonrelativistic limit, and begin by noting the electromagnetic field strength tensor:
$F^{\mu\nu}=\begin{pmatrix}0&-E_{x}&-E_{y}&-E_{z}\\ E_{x}&0&-B_{z}&B_{y}\\ E_{y}&B_{z}&0&-B_{x}\\ E_{z}&-B_{y}&B_{x}&0\end{pmatrix}$. (26)
The term $\bar{\psi}\sigma^{\mu\nu}\gamma^{5}\psi F_{\mu\nu}$ contracts to:
$2\bar{\psi}\sigma^{0i}\gamma^{5}\psi F_{0i}+\bar{\psi}\sigma^{ij}\gamma^{5}\psi F_{ij},$ (27)
where the dominant term at low velocities is the electric coupling:
$\bar{\psi}\sigma^{0i}\gamma^{5}\psi\approx\psi^{\dagger}\Sigma^{i}\psi,\qquad F_{0i}=-E^{i}.$ (28)
Thus,
$\bar{\psi}\sigma^{\mu\nu}\gamma^{5}\psi F_{\mu\nu}\to-2\,\psi^{\dagger}\vec{\Sigma}\psi\cdot\vec{E}.$ (29)
Substituting back into the Lagrangian (1), we obtain:
$\mathcal{L}_{\text{EDM}}\to d_{int}\,\psi^{\dagger}\vec{\Sigma}\psi\cdot\vec{E}=-\vec{P}(\vec{r})\cdot\vec{E},$ (30)
which corresponds to the energy density term.
Generalizing to an atomic state $\psi_{\pm}$, the total polarization may be expanded to the sum of the orbital (charge-displacement) part and the EDM (spin) part,
$\vec{P}_{\rm tot}(\vec{r})=\underbrace{\rho_{\rm charge}(\vec{r})\,\vec{r}}_{\text{orbital (Stark)}}\;+\;\underbrace{d_{int}\,\psi_{\pm}^{\dagger}(\vec{r})\,\vec{\Sigma}\,\psi_{\pm}(\vec{r})}_{\text{spin EDM}}.$ (31)
Integrating over space gives the total permanent dipole moment
$\langle\vec{d}_{\rm tot}\rangle=\int\vec{P}_{\rm tot}\,d^{3}r=\langle\vec{d}_{\rm orb}\rangle\;+\;d_{int}\,\langle\vec{\Sigma}\rangle_{\psi_{\pm}},$ (32)
$\langle\vec{d}_{\rm orb}\rangle\simeq\int\rho_{\rm charge}(\vec{r})\,\vec{r}\,d^{3}r.$ (33)
Since no electron EDM has been observed, it either vanishes or the intrinsic value $d_{\rm int}$ lies below current experimental sensitivity. Throughout this work we set $d_{\rm int}\simeq 0$, so the polarisation obtained from (31) reduces to $\vec{P}(\vec{r})\simeq\rho_{\rm charge}(\vec{r})\,\vec{r}$. To calculate orbital EDMs (e.g., the linear Stark effect), we use the effective magnetic probability current density, eq. (20), together with the polarisation equations (21) and (23). In the following we show that this relates to an effective angular-momentum like degree of freedom (a pseudo-angular momentum) that emerges when a system has an internal two- or few-level structure organised by a hidden symmetry.
IV The Stark Effect for the Hydrogen Atom
To make the analogy between magnetic and electric dipole moments concrete, we consider the classic example of the linear Stark effect in the hydrogen atom. The perturbing Hamiltonian is
$H_{S}=-\vec{d}\cdot\vec{E}=-q\vec{r}\cdot\vec{E},$ (34)
where $\vec{d}=q\vec{r}$ is the electric dipole operator and $\vec{E}$ is the applied electric field. For a field along the $z$-direction, $\vec{E}=E_{z}\hat{z}$, this becomes $H_{S}=-qzE_{z}$.
In first-order perturbation theory, the energy shift is
[…]
which vanishes unless the state has mixed parity. In this represenatation, the electron’s orbital quantum numbers are the principal quantum number $n=1,2,3,\ldots$, the orbital angular momentum $\ell=0,1,\ldots,n-1$ (with spectroscopic labels $s,p,d,f,\ldots$ for $\ell=0,1,2,3,\ldots$), and the magnetic quantum number $m=-\ell,-\ell+1,\ldots,\ell$, which set $L^{2}=\ell(\ell+1)\hbar^{2}$ and $L_{z}=m\hbar$.
In a degenerate subspace of opposite parity states, such as the $n=2$ manifold in hydrogen, the electric field couples $2s$ ($\ell=0$)and $2p$ ($\ell=1$) orbitals, leading to level splitting. Diagonalizing the perturbation in this subspace gives a permanent induced dipole moment and energy shifts linear in $E_{z}$.
This behavior mirrors the Zeeman effect, where magnetic dipole moments couple to an applied magnetic field. In both cases, an external field breaks an internal symmetry (parity or time-reversal), revealing hidden structure and inducing a dipole aligned to the field.
IV.1 Stark Effect for the $n=2$ Orbital
The hydrogen eigenstates $\psi_{n\ell m}$ separate into a radial part $R_{n\ell}(r)$ and an angular part $Y_{\ell m}(\theta,\phi)$ in spherical coordinates $(r,\theta,\phi)$ by separation of variables in the Schrödinger equation with the Coulomb potential $V(r)=-\frac{e^{2}}{4\pi\epsilon_{0}r}$. Explicitly,
$\psi_{n\ell m}(r,\theta,\phi)=R_{n\ell}(r)\,Y_{\ell m}(\theta,\phi),$ (36)
Here, we focus on the $n=2$ degenerate manifold which includes the following states: The 2s state, $\psi_{200}$, and the 2p, $\psi_{21m}$ states, with m=1,0,+1m=-1,0,+1. From (35) only the matrix element between $\psi_{200}$ and $\psi_{210}$ contributes at first order, given by,
[…]
The normalised radial wave functions given in (36) for $n=2$ are known to be;
$R_{20}(r)$ $=\frac{1}{2\sqrt{2}\,a_{0}^{3/2}}\left(2-\frac{r}{a_{0}}\right)e^{-r/(2a_{0})}$ (38)
$R_{21}(r)$ $=\frac{1}{2\sqrt{6}\,a_{0}^{3/2}}\left(\frac{r}{a_{0}}\right)e^{-r/(2a_{0})},$ (39)
and the angular wavefunctions to be;
$Y_{00}(\theta,\phi)$ $=\frac{1}{\sqrt{4\pi}},$ (40)
$Y_{10}(\theta,\phi)$ $=\sqrt{\frac{3}{4\pi}}\cos\theta$ (41)
so
$\psi_{2s}(r,\theta)$ $=R_{20}(r)\,Y_{00}=\frac{R_{20}(r)}{\sqrt{4\pi}},$ (42)
$\psi_{2p_{0}}(r,\theta)$ $=R_{21}(r)\,Y_{10}=\sqrt{\frac{3}{4\pi}}\,R_{21}(r)\cos\theta.$ (43)
The integral splits into radial and angular parts:
[…]
$\left(\int_{0}^{\infty}R_{20}(r)r^{3}R_{21}(r)\,dr\right)\cdot$
$\left(\int Y_{00}^{*}\cos\theta\,Y_{10}\,\sin\theta\,d\theta d\phi\right).$
Using the known integrals:
$\int_{0}^{\infty}R_{20}(r)\,r^{3}\,R_{21}(r)\,dr$ $=-3a_{0}$ (45)
$\int_{0}^{2\pi}d\phi\int_{0}^{\pi}Y_{00}^{*}\cos\theta\,Y_{10}\sin\theta\,d\theta$ $=1,$ (46)
we find,
[…]
Given the direction of the electric field, $\vec{E}$, defines the $z$-direction, The perturbation matrix in the $\{2s,2p_{0}\}$ basis is:
$H=\begin{pmatrix}0&3eE_{z}a_{0}\\ 3eE_{z}a_{0}&0\end{pmatrix}$ (48)
Diagonalizing this into the hybridised Stark basis,
$\psi_{2\pm}(r,\theta)=\frac{1}{\sqrt{2}}\Big[\psi_{2s}(r,\theta)\pm\psi_{2p_{0}}(r,\theta)\Big].$ (49)
yields the well known $n=2$ first-order energy Stark shifts,
$\Delta E^{(1)}_{S}=\pm 3eE_{z}a_{0}=\pm 3E_{z}d_{B}=\pm E_{z}d_{2},$ (50)
where $d_{2}$ is the magnitude of the dipole moment, with the wavefunctions given by (49), and plotted in Fig.2. A comparison of parameters for the $n=2$ and $n=3$ Stark states is shown in Tab. 1.
Figure 2: Density plot of the Stark state wavefunctions, $\psi_{2\pm}$. The asymmetry of the electron cloud around the nucleus generates an electric dipole moment.
IV.2 Dual Ohanian Model for the $n=2$ Orbital
Table 1: Calculated values for $n=2,3$, where the Stark-states have non-zero $k$ values
$n$ $g_{E}(n)=\frac{3}{2}n$ $k$ values $\int\frac{z}{r}|\psi|^{2},d^{3}r$ $\langle J_{p,z}\rangle$ $\langle d_{z}\rangle$
22 33 1,0,11,0,-1 $\tfrac{1}{2},0,\tfrac{1}{2}$ $\hbar,0,$-$\hbar$ $3d_{B},0,-3d_{B}$
33 $\tfrac{9}{2}$ 2,1,0,1,22,1,0,-1,-2 $\tfrac{2}{3},-\tfrac{1}{3},0,\tfrac{1}{3},\tfrac{2}{3}$ $2\hbar,\hbar,0,-\hbar,-2\hbar$ $\tfrac{9}{2}d_{B},\tfrac{9}{4}d_{B},0,-\tfrac{9}{4}d_{B},-\tfrac{9}{2}d_{B}$
Applying the dual Ohanian model introduced in section III.2, we calculate the effective electron charge densities from the wavefunctions,
$\rho_{2\pm}(r,\theta)=-e\,|\psi_{2\pm}(r,\theta)|^{2},$ (51)
and thus the effective microscopic polarization from the charge density,
$\vec{P}_{2\pm}(r,\theta)=r\,\rho_{2\pm}(r,\theta)\hat{r}$= (52)
$-\,\frac{e\,r}{8\pi}\Big[R_{20}^{2}+3R_{21}^{2}\cos^{2}\theta\;\pm\;2\sqrt{3}\,R_{20}R_{21}\cos\theta\Big]\hat{r}$, (53)
which is plotted in Fig. 3.
Figure 3: Effective microscopic polarization vector-fields for $n=2$ Stark states.
Figure 4: Magnetic current vector-field $\vec{J}_{m2\pm}$ calculated from (54), acting as a generator of the $\pm$ EDMs of magnitude $d_{2}$, via the left hand rule for the $n=2$ Stark states
Figure 5: The Runge-Lenz vector density, $\vec{\mathcal{A}}$, for the $n=2$ Stark states, given by (55).
Figure 6: The z-component of the Runge-Lenz vector density, $\vec{\mathcal{A}_{z}}$, for the $n=2$ Stark states, defined in (56).
Substituting (53) into (20), the effective magnetic current density may be calculated to be,
$\vec{J}_{m2\pm}=-\frac{1}{\epsilon_{0}}\nabla\times\vec{P}_{2\pm}=$ (54)
$-\frac{e\sin\theta R_{21}}{4\pi\epsilon_{0}}\left[3R_{21}(r)\cos\theta\pm\sqrt{3}R_{20}(r)\right]\hat{\phi}$
Note that $\vec{J}_{m2\pm}=\mp J_{m2}\,\hat{\phi}$ is purely azimuthal; it vanishes on the symmetry axis (where $\sin\theta=0$), and the $\pm$ sign reverses the direction between the two Stark states, as shown in Fig. 4. Furthermore, Eq. (23) implies a left-hand rule relating the magnetic current to the dipole moment, $d_{2}$. Substituting Eq. (54) into Eq. (23), one readily confirms $d_{2}=3d_{B}$.
V The Electric G-factor
The hidden Runge-Lenz (Kepler) symmetry of the Coulomb problem (see Apendix A) explains why hydrogen shows a linear Stark effect. To calculate expectation values, we can start from the Runge-Lenz density,
$\vec{\mathcal{A}}(\vec{r})=$ $\frac{\hbar^{2}}{m}\,\mathrm{Im}\!\Big[(\vec{\nabla}\psi^{\ast})\times(\vec{r}\times\vec{\nabla}\psi)\Big]-\kappa\,\frac{\vec{r}}{r}\,|\psi(\vec{r})|^{2},$ (55)
$\int\vec{\mathcal{A}}\,d^{3}r=\langle\hat{\vec{A}}\rangle$.
For the hydrogen Stark eigenstates we can choose $\psi$ real (no vector potential), so the first current-like term in (55) vanishes and the expectation value of the z-component of the Runge-Lenz vector becomes,
$\langle A_{z}\rangle=\int\mathcal{A}_{z}\,d^{3}r=-\,\kappa\int\frac{z}{r}\,|\psi(\vec{r})|^{2}\,d^{3}r,$ (56)
and the scaled vector in units of angular momentum (see appnedix A), becomes
$\langle A_{sc,z}\rangle=-\hbar n\int\frac{z}{r}\,|\psi(\vec{r})|^{2}\,d^{3}r=\langle\hat{J}_{p,z}\rangle,$ (57)
where we define $\hat{J}_{p,z}$ as the z-component pseudo angular momentum, $\hat{J}_{p}$, in parity space, which is essentially defined as the scaled Runge-Lenz vector. It follows that $\hat{J}_{p,z}$ is quantized in units of $\hbar$ so that,
$\langle J_{p,z}\rangle=\hbar\,k\ ,k=n_{1}-n_{2}\in\{-n+1,\ldots,n-1\},$ (58)
so
$\int\frac{z}{r}\,|\psi(\vec{r})|^{2}\,d^{3}r\ =\ -\,\frac{k}{n}.$ (59)
Using the analogous relation to the magnetic Landé factor defined by (7)-(8),
$\hat{\vec{d}}=\frac{g_{E}d_{B}}{\hbar}\hat{\vec{J}}_{p},$ (60)
where $g_{E}(n)=\frac{3}{2}n$ is the electric Landé factor and $d_{B}=ea_{0}$ from (15) defines the Bohr dipole scale, the expectation value of the dipole moment along the field direction and the Stark energy level shift, $\Delta E_{S}$, becomes,
$\langle d_{z}\rangle=\frac{g_{E}\,d_{B}}{\hbar}\,\langle J_{p,z}\rangle=g_{E}\,d_{B}\,k$ (61)
$\Delta E_{S}=g_{E}kd_{B}E_{z}$
These expressions show that the induced dipole moment and observed energy level shift are proportional to the projection quantum number $k$, scaled by the electric Landé factor and the fundamental dipole unit. Also, the density sum rule in (59) is a compact way to verify that the local Runge-Lenz density reproduces the global SU(2) quantization, giving a direct bridge to the electric Landé factor via (60). Just as magnetic $g$-factors differ between systems depending on their internal structure and dynamics, the electric $g$-factor encodes the efficiency with which the electron orbital in parity space is converted to a permanent polarization through the effective magnetic current loop.
Table 2: Duality between the magnetic Zeeman and electric Stark effects.
Magnetic Sector (Zeeman) Electric Sector (Linear Stark)
External field $\vec{B}$ (pseudovector) $\vec{E}$ (polar vector)
Coupled operator Total angular momentum $\hat{\vec{J}}$ Pseudo-angular momentum $\hat{\vec{J}}_{p}$
Conserved generators z-component, $\hat{J}_{z}$ z-component, Runge-Lenz vector, $\hat{A}_{sc,z}$
Bohr Model $\mu_{B}=\frac{e\hbar}{2m_{e}}$ Bohr magneton $d_{B}=\frac{2\mu_{\mathrm{B}}}{c\alpha}=ea_{0}$ Bohr EDM
Good quantum number $m_{J}$ (magnetic) $k$ (Stark or parabolic)
Dipole moment $\hat{\vec{\mu}}=g_{M}\mu_{B}\hat{\vec{J}}$$\hbar$ $\hat{\vec{d}}=g_{E}d_{B}\hat{\vec{J}}_{p}/\hbar$
Energy level Shift $\Delta E_{M}=g_{M}m_{j}\mu_{B}B_{z}$ $\Delta E_{S}=g_{E}kd_{B}E_{z}$
Hamiltonian $H_{M}=-\vec{\mu}\!\cdot\!\vec{B}$ $H_{S}=-\vec{d}\!\cdot\!\vec{E}$
Field coupling (total) $H_{M}=-(\vec{\mu}_{L}+\vec{\mu}_{S})\!\cdot\!\vec{B}$ $H_{S}=-(\vec{d}_{orb}+\vec{d}_{e})\!\cdot\!\vec{E}$
Linear response Zeeman splitting Stark splitting
Residual symmetry $SO(4)\!\to\!SO(2)$ $SO(4)\!\to\!SO(2)\!\times\!SO(2)$
Parity Symmetry Conserved (no mixing) Broken, Mixed opposite parity states
Time Reversal Sym Broken Conserved
Basis states $|\!\uparrow\rangle,|\!\downarrow\rangle$ $|+\rangle,|$-$\rangle$ (even/odd parity)
Source Circulating charge current density Circulating magnetic current density
Intrinsic operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator
Physical interpretation Precession in $\vec{B}$; magnetic polarization Polarization in $\vec{E}$; parity mixing
Geometric picture Rotation in real space Rotation in parity space
V.1 Electric Spin
To include an intrinsic dipole moment with the induced orbital effects one can combine (25) and (60) to obtain,
$\hat{\vec{d}}_{tot}=d_{B}\big[g_{E}\hat{\vec{J}}_{p}\hbar+g_{E}^{e}\,\langle\hat{\vec{S}}\rangle/\hbar\big],~~g_{E}^{e}=\frac{2d_{\rm int}}{d_{B}}$ (62)
In this case the intrinsic EDM is alined with the particles spin, and is related to a probability magnetic current produced in position space by the spin operator. Spin is a pseudovector (axial vector), it behaves like an ordinary vector under rotations, but under spatial inversion (parity) it does not change sign, unlike a true polar vector. Under time reversal it does flip sign. Consequently, an intrinsic EDM is parity-odd and time-reversal-odd, so detecting a nonzero intrinsic EDM would signal CP violation.
In contrast, there exist many fermionic quasiparticles whose two-component “spinor” resides in an internal space and couples electrically. Some canonical electric-sector realizations include graphene’s Dirac carriers, where the spinor acts on the A/B sublattice degree of freedom and couples to inversion-breaking scalar potentials introduced by substrates or gating [18, 19]; the closely related valley pseudospin associated with the two inequivalent Brillouin–zone corners, which can be driven by electric fields through inversion asymmetry and Berry-curvature (valley Hall) effects [20]; and bilayer graphene’s layer/orbital pseudospin, controllable by a perpendicular displacement field (interlayer bias) that opens and tunes an energy gap [21]. Cold-atom Fermi gases provide another electric route: two-level hyperfine manifolds act as a tunable pseudospin-$\tfrac{1}{2}$ coherently manipulated via optical (electric-dipole) Raman coupling and AC–Stark shifts [22].
In this sense, the Stark/EDM pseudo-angular momentum $\hat{\vec{J}}_{p}$ is of the same kind: an orbital, internal $SU(2)$ (Runge-Lenz) that couples opposite-parity components within a fixed $n$ manifold, while the underlying particles remain bona fide fermions.
VI Discussion and Implications
The framework presented here provides a concrete route to visualizing electric dipole moments in analogy to magnetic dipole moments, with a comparison shown in Tab. 2. By introducing the concept of an effective magnetic current associated with permanent polarization, we present a dual description of a permanent EDM and motivate the definition of an electric Bohr moment and electric $g$-factor. This duality between magnetic and electric dipole systems enables deeper insight into the symmetry properties of the underlying fields. It highlights how internal circulations, whether of electric or magnetic nature, can result in observable moments coupled to external fields, and how these couplings break fundamental symmetries. Furthermore, the quantized boundary current model connects well with modern polarization theory in materials, linking quantum Berry phase descriptions with classical electromagnetic analogies.
By analogy, classical polarized electric dipoles are naturally described using the impressed-source and field-equivalence formalisms [23, 24]. The driving EMF that separates charge is represented as an external impressed electric field, equivalently as a polarisation field with nonzero curl. Because the EMF is localised, the boundary carries (i) dipolar charge that terminates the polarisation and (ii) a fictitious magnetic surface current parallel to the polarisation. This description captures the external fields and terminal responses without modelling the internal workings of the (static or dynamic) source. This viewpoint underpins our recent analyses of active classical dipoles, which employ such magnetic equivalence currents [12, 13].
The Stark–sector theory developed here is the quantum dual of that classical picture. Starting from the wavefunction’s microscopic polarization $\vec{P}(\vec{r})$, a nonzero curl defines an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$, in a similar way to the classical picture [12, 13, 14, 25]. Thus, the induced EDM in a static field is governed by a circulating magnetic probability current in the orbital (Runge-Lenz) sector, mirroring how an impressed electric field in classical circuit and antenna theory is traded for an equivalent magnetic surface current that generates the same exterior fields. In this sense the dual-Ohanian formulation makes explicit the classical-quantum correspondence, whereby an impressed electric drive is represented by a magnetic equivalence current in the classical dipole [12, 13], while Stark mixing is represented by a magnetic probability current in the quantum dipole.
Appendices
Appendix A Runge-Lenz Operator
The unperturbed hydrogen atom is described by the Coulomb Hamiltonian
$\hat{H}_{0}=\frac{\hat{\vec{p}}^{2}}{2\mu}-\frac{\kappa}{\hat{r}},\qquad\mu=\frac{m_{e}m_{p}}{m_{e}+m_{p}}\ \ [\mathrm{kg}],$ (63)
where $\mu$ is the reduced mass of the electron-proton system. The Coulomb coupling constant is given by
$\kappa=\frac{e^{2}}{4\pi\epsilon_{0}}\ \ [\mathrm{J\,m}],\qquad e\ \ [\mathrm{C}],\quad\epsilon_{0}\ \ [\mathrm{F\,m^{-1}}],$ (64)
and the characteristic Bohr radius follows as
$a_{0}=\frac{4\pi\epsilon_{0}\hbar^{2}}{\mu e^{2}}\ \ [\mathrm{m}],\qquad\hbar\ \ [\mathrm{J\,s}].$ (65)
The canonical position and momentum operators are defined as
$\hat{\vec{r}}=(\hat{x},\hat{y},\hat{z}),\quad\hat{r}=|\hat{\vec{r}}|,\quad\widehat{\vec{r}}=\frac{\hat{\vec{r}}}{\hat{r}}$, (66)
$\hat{\vec{p}}=-\,i\hbar\,\vec{\nabla}\ \ [\mathrm{kg\,m\,s^{-1}}],$ (67)
and the orbital angular momentum operator takes the standard form
$\hat{\vec{L}}=\hat{\vec{r}}\times\hat{\vec{p}}\ \ [\mathrm{J\,s}].$ (68)
A less familiar but equally fundamental conserved quantity of the Coulomb problem is the Runge-Lenz vector [26, 27, 28, 29], defined (in symmetrized operator form) as
$\hat{\vec{A}}=\frac{1}{2\mu}\big(\hat{\vec{p}}\times\hat{\vec{L}}-\hat{\vec{L}}\times\hat{\vec{p}}\big)-\kappa\,\frac{\widehat{\vec{r}}}{r}\ \ [\mathrm{J\,m}].$ (69)
This operator generates a hidden dynamical symmetry of the hydrogen atom, corresponding to the conserved direction of the elliptical orbit in the classical limit. Its magnitude and commutation relations with $\hat{\vec{L}}$ lead to the well-known $O(4)\!\simeq\!SU(2)\!\times\!SU(2)$ symmetry of the bound-state spectrum [26].
To obtain an angular-momentum-dimension, we define a scaled Runge-Lenz operator,
$\hat{\vec{A}}_{\mathrm{sc}}\equiv\frac{\sqrt{\mu}\hat{\vec{A}}}{\sqrt{-2\hat{H}_{0}}}\ \ [\mathrm{J\,s}],\qquad\hat{\vec{J}}_{p}\equiv\hat{\vec{A}}_{\mathrm{sc}},$ (70)
which introduces a pseudo-angular momentum $\hat{\vec{J}}_{p}$ with the same algebraic structure as $\hat{\vec{L}}$.
On a hydrogenic energy eigenspace $\hat{H}_{0}|_{E_{n}}=E_{n}|_{E_{n}}$ with
$E_{n}=-\,\frac{\mu\kappa^{2}}{2\hbar^{2}n^{2}},$ (71)
the scaling factor becomes a number:
$\sqrt{-2E_{n}/\mu}=\frac{\kappa}{\hbar n}\Rightarrow\hat{\vec{A}}_{\mathrm{sc}}=\frac{\hbar n}{\kappa}\,\hat{\vec{A}}=\frac{n}{c\alpha}\,\hat{\vec{A}}$ (72)
on the $n$ manifold. Thus, within each principal quantum manifold $n$, the scaled Runge-Lenz vector plays the role of a conserved pseudo-angular momentum, forming with $\hat{\vec{L}}$ the generators of the $SU(2)\times SU(2)$ symmetry that underlies the degeneracy of the hydrogen spectrum.
Appendix B Hydrogen quantum numbers (spherical vs. parabolic)
In hydrogen, the electron’s orbital quantum numbers are the principal quantum number $n=1,2,3,\ldots$, the orbital angular momentum $\ell=0,1,\ldots,n-1$ (with spectroscopic labels $s,p,d,f,\ldots$ for $\ell=0,1,2,3,\ldots$), and the magnetic quantum number $m=-\ell,-\ell+1,\ldots,\ell$, which set $L^{2}=\ell(\ell+1)\hbar^{2}$ and $L_{z}=m\hbar$. In an external electric field it is convenient to use the parabolic set $(n_{1},n_{2},m)$, obtained by separating the Schrödinger equation in parabolic coordinates $(\xi=r+z,\ \eta=r-z,\ \phi)$. Here $n_{1},n_{2}=0,1,2,\ldots$ are not principal quantum numbers: they count radial-like nodes along the two parabolic directions and, together with $m$, obey the constraint $n_{1}+n_{2}+|m|+1=n$ (where $n$ remains the single principal quantum number). The combination $k\equiv n_{1}-n_{2}\in\{-(n-1),\ldots,n-1\}$ labels the linear Stark splitting and controls the expectation values $\langle z\rangle=\tfrac{3}{2}\,n\,a_{0}\,k$ and $\langle A_{\mathrm{sc},z}\rangle=\hbar k$. In zero field, the parabolic basis is just another unitary choice within the $n$ manifold, equivalent to the spherical $\lvert n\ell m\rangle$ basis.
Acknowledgements.
This work was funded by the ARC Centre of Excellence for Dark Matter Particle Physics, CE200100008.

### v2

@@
 Here the spin operator, $\hat{\vec{S}}$, encodes any intrinsic EDM, $d_{\rm int}$, from symmetry-violating interactions, while $\hat{\vec{J}}_{p}$ captures the Stark-induced pseudo-angular momentum set by Runge-Lenz symmetry.
-This dual framework shows that induced EDMs arise from effective circulating probability magnetic currents, directly mirroring magnetic dipole moments arising from circulating probability electric currents.
+In this dual framework, an induced EDM, which physically originates from charge displacement via parity mixing, can be equivalently represented by an effective circulating probability magnetic current, mirroring the probability electric-current representation of magnetic dipole moments.
 I Introduction
@@
 II Magnetic Dipole Moments
-II.1 The Semi-Classical Bohr Magnetron
+II.1 The Semi-Classical Bohr Magneton
 The Bohr magneton arises naturally from a semi-classical model of an electron orbiting a nucleus, as in the Bohr model of the atom.
@@
 In the Standard Model, such EDMs are predicted to exist but at levels far below current experimental sensitivity (for example, the electron EDM is expected to be about $10^{-38}\,\mathrm{e\cdot cm}$, whereas the neutron is about $10^{-31}\,\mathrm{e\cdot cm}$).
-State-of-the-art experiments constrain the electron EDM to be smaller than $1.3\times 10^{-29}\,\mathrm{e\cdot cm}$ and the neutron EDM below $1.0\times 10^{-26}\,\mathrm{e\cdot cm}$, with similar null results for the muon and proton.
+State-of-the-art experiments constrain the electron EDM to be smaller than $4.1\times 10^{-30}\,\mathrm{e\cdot cm}$ [18] and the neutron EDM below $1.8\times 10^{-26}\,\mathrm{e\cdot cm}$ [19], with similar null results for the muon and proton.
 These increasingly tight bounds are crucial because a confirmed intrinsic EDM would signal new CP-violating physics beyond the Standard Model, possibly helping to explain the matter-antimatter asymmetry of the Universe.
@@
 In this case, the dipole operator acts on a pseudo-angular momentum degree of freedom, defined in the two-level parity subspace, and the resulting EDM depends on the external field and the mixing angle.
-III.1 Semi-Classical Model of a Quanta of Permanent Polarization and dipole moment
-Figure 1: Semi-classical representations of the Bohr EDM, $\vec{d}_{B}$.
-(a) Ideal semi-classical unit cell of quantised constant permanent polarization, $\Delta\vec{P}=\frac{e}{\pi a_{e}^{2}}\hat{z}$.
-In this model, the total angular momentum, and net electrical and magnetic charge are zero.
-However, because the curl of the polarization is non-zero, an effective magnetic current appears on the radial surface, which can be modelled as a quanta of a south and north pole ($\mp q_{m}=\mp\frac{h}{e}$) rotating in opposite directions, and is related to pseudo-spin (or pseudo-angular momentum) in parity space rather than position space.
-(b) Classical Bohr EDM, represented by two point charges $\pm e$ separated by a distance of the Bohr radius.
-(c) Schematic representing a semi-classical intrinsic Bohr EDM, connected to the position space spin vector, $\vec{S}$, related to an effective magnetic current of $\vec{I}_{m}$ at a radial distance of $a_{0}$.
-Similar to the semi-classical atomic Bohr magneton calculation one can use a similar technique to derive the semi-classical atomic dual electrical analogue.
-To do this we conceive a toy model model of two quanta of opposite signed magnetic charge $\pm\frac{h}{e}$ as shown in Fig.1, rotating at the the electron Bohr radius, $a_{0}=\frac{\hbar}{m_{\mathrm{e}}c\alpha}$, creating a EDM.
-Here $\alpha=\frac{e^{2}}{4\pi\varepsilon_{0}\hbar c}$ is the fine structure constant, $\epsilon_{0}$ the permittivity of free space, $c$ the speed of light.
-In this model, the total net electrical and magnetic charge as well as angular momentum is zero.
-However a finite effective magnetic current exits, which we pose is not generated by mechanical orbital motion, but by pseudospin parity mixing.
-Given that $q_{m}=\frac{h}{e}$, the magnetic current at the boundary is given by,
-$\vec{I}_{m}=-\frac{2q_{m}v}{2\pi a_{0}}\hat{\phi}=-\frac{e}{2\pi a_{0}^{2}m^{*}\epsilon_{0}\alpha c}(a_{0}m^{*}v)\hat{\phi}$ (14)
-Here $m^{*}$ is the effective mass of the atomic particle related to the permanent dipole, which can vary from $0.01m_{e}$ to very large.
-Assuming the total mass per unit cell (Fig.1(c)) is equivalent to the effective mass of the electron, then we assume $m^{*}\sim\frac{m_{e}}{2}$ (two magnetic charged particles), and equating $a_{0}m^{*}v\sim\hbar$, as in the derivation of the Bohr magneton in (5), then the magnetic current becomes, $\vec{I}_{m}=-\frac{e\hbar}{\pi a_{0}^{2}m_{e}\epsilon_{0}\alpha c}\hat{\phi}=-\frac{2\mu_{\mathrm{B}}}{\pi a_{0}^{2}\epsilon_{0}\alpha c}\hat{\phi}$.
-Thus, we can define the “Bohr EDM” in terms of the effective azimuthal magnetic current, via a left hand rule,
+III.1 Semi-Classical Mnemonic for a Quantum of Permanent Polarization and Dipole Moment
+Figure 1: Semi-classical representations of the Bohr EDM scale, $\vec{d}_{B}$.
+(a) Idealized quasi-1D “unit cell” of quantized permanent polarization, $\Delta\vec{P}=\frac{e}{\pi a_{0}^{2}}\hat{z}$, for which $\nabla\times\vec{P}\neq 0$ implies an effective magnetic current on the radial surface (a dual representation of the polarization texture, not a literal magnetic-charge flow).
+For mnemonic purposes this surface current can be pictured as two fictitious opposite magnetic charges $\mp q_{m}=\mp\frac{h}{e}$ circulating in opposite senses, associated with pseudo-spin (pseudo-angular momentum) in parity space rather than mechanical orbital motion.
+(b) Classical Bohr EDM: two point charges $\pm e$ separated by the Bohr radius $a_{0}$.
+(c) Schematic of an intrinsic Bohr-scale EDM connected to the position-space spin vector $\vec{S}$, represented equivalently by an effective azimuthal magnetic current $\vec{I}_{m}$ at radius $\sim a_{0}$.
+We define the Bohr EDM scale as
+$d_{B}\;\equiv\;ea_{0},$ (14)
+which is the natural atomic dipole unit.
+In the Coulomb problem this scale is consistent with the hydrogenic $SO(4)$ (Runge–Lenz) structure within a fixed $n$-manifold, and it also has an immediate classical interpretation: it is the dipole moment of two point charges $\pm e$ separated by a distance $a_{0}$ [Fig.
+1(b)].
+Furthermore, the same dipole unit appears in a quasi-1D “cell” formulation governed by modern polarization theory, when the size of the cell spacing is of order $a_{0}$.
+The following toy model is included only as a heuristic mnemonic (not as a physical derivation) to connect this dipole scale to a dual-current picture.
+Similar to the semi-classical Bohr magneton estimate, we imagine an effective surface magnetic current associated with the curl of the polarization.
+For mnemonic purposes, this may be visualized as two fictitious opposite “magnetic charges” $\pm\frac{h}{e}$ circulating at the Bohr radius $a_{0}=\frac{\hbar}{m_{\mathrm{e}}c\alpha}$, where $\alpha=\frac{e^{2}}{4\pi\varepsilon_{0}\hbar c}$.
+In this picture, the net electric charge, net magnetic charge, and total angular momentum of the cell are zero, but an effective magnetic surface current exists, which we associate with pseudo-spin (parity mixing) rather than mechanical orbital motion.
+With $q_{m}=\frac{h}{e}$, the effective magnetic current at the boundary is written as
+$\vec{I}_{m}=-\frac{2q_{m}v}{2\pi a_{0}}\hat{\phi}=-\frac{e}{2\pi a_{0}^{2}m^{*}\epsilon_{0}\alpha c}(a_{0}m^{*}v)\hat{\phi}$, (15)
+where $m^{*}$ is an effective inertia parameter for the internal (parity-space) degree of freedom.
+For a Bohr-scale dimensional estimate we take the usual quantization scale $a_{0}m^{*}v\sim\hbar$ with $m^{*}\sim\mathcal{O}(m_{e})$ (the precise value of $m^{*}$ is not used elsewhere in the manuscript).
+This gives
+$\vec{I}_{m}\approx-\frac{e\hbar}{\pi a_{0}^{2}m_{e}\epsilon_{0}\alpha c}\hat{\phi}=-\frac{2\mu_{\mathrm{B}}}{\pi a_{0}^{2}\epsilon_{0}\alpha c}\hat{\phi}$, (16)
+and we may express the Bohr EDM in terms of the effective azimuthal magnetic current (left-hand rule) as
 $d_{B}=-\epsilon_{0}I_{m}\pi a_{0}^{2}=\frac{2\mu_{\mathrm{B}}}{c\alpha}=ea_{0},$ (17)
-the dual of the Bohr magneton in (5).
-This dipole moment is equivalent to two charges $\pm e$ separated by a distance of $a_{0}$ as shown in Fig.
-1 (b), which also means $\vec{a}=a_{0}\hat{z}$, so Fig.
-1(a) and 1(b) are equivalent.
+the electromagnetic dual of the Bohr magneton in (5). Thus Fig. 1(a) and Fig. 1(b) represent the same Bohr-scale dipole unit: $\vec{a}=a_{0}\hat{z}$ implies $d_{B}=ea_{0}$.
 In modern polarization theory, the bulk polarization is defined “modulo a quantum”, along with a Berry phase defined modulo $2\pi$, and is given by
@@
 This follows from an electron in a periodic potential, which gives the $U(1)$ gauge freedom (Bloch’s theorem); $\psi_{n\vec{k}}\!\to\!e^{-i\vec{k}\!\cdot\!\vec{a}}\psi_{n\vec{k}}$, which shifts the Berry phase by $2\pi$ and $\vec{P}$ by $e\vec{a}/V$.
-Here we consider a quasi-1D “cell” as shown in Fig.1 (c) of length $a$ (along $\hat{\vec{z}}$) and cross-section $A_{\!\mathrm{eff}}$, so the effective volume is, $V_{\!\mathrm{eff}}\;=\;A_{\!\mathrm{eff}}\,a$.
+Here we consider a quasi-1D “cell” as shown in Fig.
+1 of length $a$ (along $\hat{\vec{z}}$) and cross-section $A_{\!\mathrm{eff}}$, so the effective volume is $V_{\!\mathrm{eff}}=A_{\!\mathrm{eff}}a$.
 In our construction $2\pi$ winding of the internal phase is equivalent to choosing $\vec{a}=a\hat{\vec{z}}$.
@@
 $\Delta P_{z}\;=\;\frac{e\,a}{V_{\!\mathrm{eff}}}\;=\;\frac{e}{A_{\!\mathrm{eff}}}\;=\;\frac{e}{\pi a_{0}^{2}}.$ (19)
-This is equivalent to a single cell of a super cell in modern polarization theory [10], with a quanta of permanent polarization, $\Delta P_{z}$ of size on the order of the Bohr radius, where $A_{\!\mathrm{eff}}=\pi a_{0}^{2}$, as shown in Fig.1.
-Asserting that $\vec{a}=a_{0}\hat{z}$, the dipole moment can then be calculated from $\vec{d}_{B}=\int\Delta P_{z}d^{3}r=ea_{0}\hat{z}=d_{B}\hat{z}$, consistent with (15).
+This is equivalent to a single cell of a supercell in modern polarization theory [10], with a quantum of permanent polarization $\Delta P_{z}$ set by $A_{\!\mathrm{eff}}=\pi a_{0}^{2}$, as shown in Fig.
+1.
+With $\vec{a}=a_{0}\hat{z}$, the dipole moment follows as $\vec{d}_{B}=\int\Delta P_{z}\,d^{3}r=ea_{0}\hat{z}=d_{B}\hat{z}$, consistent with (17).
+For completeness when discussing the hydrogenic problem, we note that, for a finite-mass nucleus, the mass parameter in the internal (relative-coordinate) Coulomb problem should be understood as the reduced mass
+$\mu\;=\;\frac{m_{e}m_{N}}{m_{e}+m_{N}},$ (20)
+rather than the electron mass $m_{e}$, so that the Bohr radius is replaced by
+$a_{\mu}\;=\;\frac{4\pi\epsilon_{0}\hbar^{2}}{\mu Ze^{2}},$ (21)
+and hence the Bohr dipole scale used throughout becomes $d_{B}=ea_{\mu}$ (rather than $ea_{0}$ defined using $m_{e}$).
+This substitution rescales lengths and energies but does not modify the underlying symmetry structure: the $SO(4)\simeq SU(2)\times SU(2)$ algebra and associated commutation relations remain unchanged (see Appendix C), with $m\rightarrow\mu$ in the Hamiltonian.
+In practice, for hydrogen the correction is very small,
+$\frac{\mu}{m_{e}}\;=\;\frac{1}{1+m_{e}/m_{p}}\;\approx\;0.99946,$ (22)
+so $a_{\mu}$ (and thus $d_{B}$) differs from the $m_{e}$ value by only $\sim 5.4\times 10^{-4}$ (about $0.054\%$), and either convention may be adopted depending on the desired level of precision.
 III.2 Dual Ohanian Model: Dipole moment from Internal Polarisation
@@
 $\vec{P}(\vec{r})$ $=d_{B}\,\psi^{\dagger}(\vec{r})\,\vec{\Sigma}\,\psi(\vec{r}),\qquad d_{B}\equiv ea_{0},$ (23)
-$=\frac{2d_{B}}{\hbar}\psi^{\dagger}(\vec{r})\hat{\vec{S}}\psi(\vec{r}),\qquad\hat{\vec{S}}\equiv\frac{\hbar}{2}\bm{\sigma}.$ (19)
-where, $\bm{\sigma}$ denotes the vector of Pauli operators acting on the two-level parity space, $\vec{\Sigma}$ is the Pauli spin matrix vector and $\psi$ the Dirac or Pauli spinor.
+$=\frac{2d_{B}}{\hbar}\psi^{\dagger}(\vec{r})\hat{\vec{S}}\psi(\vec{r}),\qquad\hat{\vec{S}}\equiv\frac{\hbar}{2}\vec{\sigma}.$ (24)
+where, $\vec{\sigma}$ denotes the vector of Pauli operators acting on the two-level parity space, $\vec{\Sigma}$ is the Pauli spin matrix vector and $\psi$ the Dirac or Pauli spinor.
 Adopting the dual (Weber) convention for magnetic current, the curl of $\vec{P}$ defines an effective probability magnetic current density,
@@
 If the electron carries an intrinsic electric dipole moment, $d_{int}$, the local microscopic polarization density of the electron wavefunction would be of the form,
-$\vec{P}_{\rm e}(\vec{r})=d_{int}\,\psi^{\dagger}(\vec{r})\,\bm{\Sigma}\,\psi(\vec{r})\;\;\approx\;\;\frac{2d_{int}}{\hbar}\,\vec{s}(\vec{r}),$ (24)
+$\vec{P}_{\rm e}(\vec{r})=d_{int}\,\psi^{\dagger}(\vec{r})\,\vec{\Sigma}\,\psi(\vec{r})\;\;\approx\;\;\frac{2d_{int}}{\hbar}\,\vec{s}(\vec{r}),$ (29)
 consistent with
 $\hat{\vec{d}}_{e}=\frac{g_{E}^{e}d_{B}}{\hbar}\hat{\vec{S}}~~~~g_{E}^{e}=\frac{2d_{int}}{d_{B}},$ (30)
-where $\vec{s}(\vec{r})=\psi^{\dagger}(\vec{r})\frac{\hbar}{2}\bm{\sigma}\psi(\vec{r})$ is the nonrelativistic spin density and $g_{E}^{e}$ is the effective electric Landé g-factor of the electron.
+where $\vec{s}(\vec{r})=\psi^{\dagger}(\vec{r})\frac{\hbar}{2}\vec{\sigma}\psi(\vec{r})$ is the nonrelativistic spin density and $g_{E}^{e}$ is the effective electric Landé g-factor of the electron.
 To understand how the electron behaves in an electric field we consider the interaction term in (1).
@@
 Since no electron EDM has been observed, it either vanishes or the intrinsic value $d_{\rm int}$ lies below current experimental sensitivity.
-Throughout this work we set $d_{\rm int}\simeq 0$, so the polarisation obtained from (31) reduces to $\vec{P}(\vec{r})\simeq\rho_{\rm charge}(\vec{r})\,\vec{r}$.
+Throughout this work we set $d_{\rm int}\simeq 0$, so that the polarization obtained from (36) reduces to $\vec{P}(\vec{r})\simeq\rho_{\rm charge}(\vec{r})\,\vec{r}$.
 To calculate orbital EDMs (e.g., the linear Stark effect), we use the effective magnetic probability current density, Eq.
-(20), together with the polarisation equations (21) and (23).
-In the following we show that this relates to an effective angular-momentum like degree of freedom (a pseudo-angular momentum) that emerges when a system has an internal two- or few-level structure organised by a hidden symmetry.
+(25), together with the polarization relations (26) and (28).
+Importantly, the induced orbital EDM is produced physically by parity mixing (charge displacement); the associated magnetic-current description is a topological field equivalence (a dual representation of the polarization texture), not an additional causal mechanism.
+We further show that this charge displacement can be organized in terms of an effective angular-momentum like degree of freedom, a pseudo-angular momentum in parity space, that emerges when the system possesses an internal two- (or few-) level structure governed by a hidden symmetry.
 IV The Stark Effect for the Hydrogen Atom
@@
 33 $\tfrac{9}{2}$ 2,1,0,1,22,1,0,-1,-2 $\tfrac{2}{3},-\tfrac{1}{3},0,\tfrac{1}{3},\tfrac{2}{3}$ $2\hbar,\hbar,0,-\hbar,-2\hbar$ $\tfrac{9}{2}d_{B},\tfrac{9}{4}d_{B},0,-\tfrac{9}{4}d_{B},-\tfrac{9}{2}d_{B}$
-Applying the dual Ohanian model introduced in section III.2, we calculate the effective electron charge densities from the wavefunctions,
+Applying the dual Ohanian model introduced in section III.2, we calculate the effective electron charge densities from the wavefunctions.
+Note, here the electron coordinates are relative to the nucleus and that the associated microscopic “effective” fields are internal fields defined in the atomic rest frame,
 $\rho_{2\pm}(r,\theta)=-e\,|\psi_{2\pm}(r,\theta)|^{2},$ (56)
@@
 Source Circulating charge current density Circulating magnetic current density
-Intrinsic operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator
+Intrinsic operator $\hat{\vec{S}}=\frac{\hbar}{2}\vec{\sigma}$ Spin operator $\hat{\vec{S}}=\frac{\hbar}{2}\vec{\sigma}$ Spin operator
 Physical interpretation Precession in $\vec{B}$; magnetic polarization Polarization in $\vec{E}$; parity mixing
@@
 This viewpoint underpins our recent analyses of active classical dipoles, which employ such magnetic equivalence currents [12, 13].
-The Stark–sector theory developed here is the quantum dual of that classical picture.
-Starting from the wavefunction’s microscopic polarization $\vec{P}(\vec{r})$, a nonzero curl defines an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$, in a similar way to the classical picture [12, 13, 14, 25].
-Thus, the induced EDM in a static field is governed by a circulating magnetic probability current in the orbital (Runge-Lenz) sector, mirroring how an impressed electric field in classical circuit and antenna theory is traded for an equivalent magnetic surface current that generates the same exterior fields.
-In this sense the dual-Ohanian formulation makes explicit the classical-quantum correspondence, whereby an impressed electric drive is represented by a magnetic equivalence current in the classical dipole [12, 13], while Stark mixing is represented by a magnetic probability current in the quantum dipole.
+The Stark-sector theory developed here is the quantum dual of the classical equivalence-current picture.
+Starting from the wavefunction’s microscopic polarization $\vec{P}(\vec{r})$, we define an effective magnetic probability current, $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$, which provides a dual-current representation of the polarization texture in close analogy with the classical construction [12, 13, 14, 27].
+Importantly, the induced EDM in a static field arises physically from charge displacement via parity (Stark) mixing; $\vec{J}_{m}$ is not an additional causal mechanism, but a convenient field-equivalent description that encodes the same orbital (Runge-Lenz) structure.
+This mirrors how, in classical circuit and antenna theory, an impressed electric drive can be traded for an equivalent magnetic surface current that reproduces the same exterior fields.
+In this sense, the dual-Ohanian formulation makes the classical-quantum correspondence explicit: an impressed electric excitation is represented by a magnetic equivalence current in the classical dipole [12, 13], while Stark mixing is represented by the magnetic probability-current representation of the quantum dipole.
 Appendices
@@
 In zero field, the parabolic basis is just another unitary choice within the $n$ manifold, equivalent to the spherical $\lvert n\ell m\rangle$ basis.
+Appendix C Hydrogenic $SO(4)$ structure and the pseudo-angular momentum
+For a hydrogenic atom/ion, the following familiar Coulomb-problem commutation relations encode the hidden $SO(4)$ symmetry of the bound-state ($H<0$) spectrum.
+$[L_{i},L_{j}]$ $=i\hbar\,\epsilon_{ijk}L_{k},$ (78)
+$[L_{i},A_{j}]$ $=i\hbar\,\epsilon_{ijk}A_{k},$ (79)
+$[A_{i},A_{j}]$ $=-2\mu H\,i\hbar\,\epsilon_{ijk}L_{k},$ (80)
+For bound states it is convenient to use the scaled Runge–Lenz operator (75) , so (80) implies
+$A_{{\rm sc},i},A_{{\rm sc},j}]\;=\;i\hbar\,\epsilon_{ijk}L_{k},$ (81)
+and $\vec{A}_{\rm sc}$ has the dimensions of angular momentum (like $\vec{L}$). One may then form the standard $SO(4)\simeq SU(2)\times SU(2)$ generators
+$\vec{J}_{\pm}\;\equiv\;\frac{1}{2}\left(\vec{L}\pm\vec{A}_{\rm sc}\right),$ (82)
+which satisfy the $SU(2)$ algebras
+$[J_{\pm,i},J_{\pm,j}]$ $=i\hbar\,\epsilon_{ijk}J_{\pm,k},$ (83)
+$[J_{+,i},J_{-,j}]$ $=0.$ (84)
+Within a fixed $n$-manifold (degenerate subspace), these relations provide the rigorous $SU(2)$ structure underpinning the Landé-factor analogy used below.
+In the Stark/parity (two- or few-level) subspace relevant for the orbital EDM, we identify the pseudo-angular momentum $\vec{J}_{p}$ with the appropriate projected $SU(2)$ generator (equivalently, the corresponding component of $\vec{J}_{\pm}$ along the field axis) that labels the parabolic/Stark structure.
+Consequently, $\vec{J}_{p}$ obeys the standard $SU(2)$ commutation relations,
+$J_{p,i},J_{p,j}]\;=\;i\hbar\,\epsilon_{ijk}J_{p,k},$ (85)
+justifying the Landé-factor form $\hat{\vec{d}}_{\rm orb}=g_{E}d_{B}\,\hat{\vec{J}}_{p}/\hbar$ used in this work.
 Acknowledgements.